\begin{thebibliography}{63}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Abid et~al.(2021)Abid, Farooqi, and Zou}]{abid2021persistent}
Abubakar Abid, Maheen Farooqi, and James Zou. 2021.
\newblock Persistent anti-muslim bias in large language models.
\newblock In \emph{Proceedings of the 2021 AAAI/ACM Conference on AI, Ethics, and Society}, pages 298--306.

\bibitem[{Agarwal et~al.(2018)Agarwal, Beygelzimer, Dud{\'\i}k, Langford, and Wallach}]{agarwal2018reductions}
Alekh Agarwal, Alina Beygelzimer, Miroslav Dud{\'\i}k, John Langford, and Hanna Wallach. 2018.
\newblock A reductions approach to fair classification.
\newblock In \emph{International conference on machine learning}, pages 60--69. PMLR.

\bibitem[{Anahideh et~al.(2022)Anahideh, Asudeh, and Thirumuruganathan}]{anahideh2022fair}
Hadis Anahideh, Abolfazl Asudeh, and Saravanan Thirumuruganathan. 2022.
\newblock Fair active learning.
\newblock \emph{Expert Systems with Applications}, 199:116981.

\bibitem[{Asad et~al.(2019)Asad, Dombrowski, Costanza-Chock, Erete, and Harrington}]{asad2019academic}
Mariam Asad, Lynn Dombrowski, Sasha Costanza-Chock, Sheena Erete, and Christina Harrington. 2019.
\newblock Academic accomplices: Practical strategies for research justice.
\newblock In \emph{Companion Publication of the 2019 on Designing Interactive Systems Conference 2019 Companion}, pages 353--356.

\bibitem[{Balayn et~al.(2018)Balayn, Mavridis, Bozzon, Timmermans, and Szl{\'a}vik}]{balayn2018characterising}
Agathe Balayn, Panagiotis Mavridis, Alessandro Bozzon, Benjamin Timmermans, and Zolt{\'a}n Szl{\'a}vik. 2018.
\newblock Characterising and mitigating aggregationbias in crowdsourced toxicity annotations.
\newblock In \emph{Proceedings of the 1st Workshop on Disentangling the Relation between Crowdsourcing and Bias Management}. CEUR.

\bibitem[{Bender et~al.(2021)Bender, Gebru, McMillan-Major, and Shmitchell}]{bender2021dangers}
Emily~M Bender, Timnit Gebru, Angelina McMillan-Major, and Shmargaret Shmitchell. 2021.
\newblock On the dangers of stochastic parrots: Can language models be too big?
\newblock In \emph{Proceedings of the 2021 ACM conference on fairness, accountability, and transparency}, pages 610--623.

\bibitem[{Bless and Fiedler(2014)}]{bless2014social}
Herbert Bless and Klaus Fiedler. 2014.
\newblock \emph{Social cognition: How individuals construct social reality}.
\newblock Psychology Press.

\bibitem[{Blodgett et~al.(2020)Blodgett, Barocas, Daum{\'e}~III, and Wallach}]{blodgett2020language}
Su~Lin Blodgett, Solon Barocas, Hal Daum{\'e}~III, and Hanna Wallach. 2020.
\newblock Language (technology) is power: A critical survey of" bias" in nlp.
\newblock \emph{arXiv preprint arXiv:2005.14050}.

\bibitem[{Bojke et~al.(2021)Bojke, Soares, Claxton, Colson, Fox, Jackson, Jankovic, Morton, Sharples, and Taylor}]{bojke2021developing}
Laura Bojke, Marta Soares, Karl Claxton, Abigail Colson, Aim{\'e}e Fox, Christopher Jackson, Dina Jankovic, Alec Morton, Linda Sharples, and Andrea Taylor. 2021.
\newblock Developing a reference protocol for structured expert elicitation in health-care decision-making: a mixed-methods study.
\newblock \emph{Health Technology Assessment (Winchester, England)}, 25(37):1.

\bibitem[{Bolukbasi et~al.(2016)Bolukbasi, Chang, Zou, Saligrama, and Kalai}]{bolukbasi2016man}
Tolga Bolukbasi, Kai-Wei Chang, James~Y Zou, Venkatesh Saligrama, and Adam~T Kalai. 2016.
\newblock Man is to computer programmer as woman is to homemaker? debiasing word embeddings.
\newblock \emph{Advances in neural information processing systems}, 29.

\bibitem[{Cabrera et~al.(2014)Cabrera, Miller, and Schneider}]{cabrera2014systematic}
Guillermo~F Cabrera, Christopher~J Miller, and Jeff Schneider. 2014.
\newblock Systematic labeling bias: De-biasing where everyone is wrong.
\newblock In \emph{2014 22nd International Conference on Pattern Recognition}, pages 4417--4422. IEEE.

\bibitem[{Calmon et~al.(2017)Calmon, Wei, Vinzamuri, Natesan~Ramamurthy, and Varshney}]{calmon2017optimized}
Flavio Calmon, Dennis Wei, Bhanukiran Vinzamuri, Karthikeyan Natesan~Ramamurthy, and Kush~R Varshney. 2017.
\newblock Optimized pre-processing for discrimination prevention.
\newblock \emph{Advances in neural information processing systems}, 30.

\bibitem[{Crawford(2016)}]{crawford2016artificial}
Kate Crawford. 2016.
\newblock Artificial intelligence’s white guy problem.
\newblock \emph{The New York Times}, 25(06):5.

\bibitem[{D{\'\i}az et~al.(2022)D{\'\i}az, Kivlichan, Rosen, Baker, Amironesei, Prabhakaran, and Denton}]{diaz2022crowdworksheets}
Mark D{\'\i}az, Ian Kivlichan, Rachel Rosen, Dylan Baker, Razvan Amironesei, Vinodkumar Prabhakaran, and Emily Denton. 2022.
\newblock Crowdworksheets: Accounting for individual and collective identities underlying crowdsourced dataset annotation.
\newblock In \emph{Proceedings of the 2022 ACM Conference on Fairness, Accountability, and Transparency}, pages 2342--2351.

\bibitem[{Dwork et~al.(2012)Dwork, Hardt, Pitassi, Reingold, and Zemel}]{dwork2012fairness}
Cynthia Dwork, Moritz Hardt, Toniann Pitassi, Omer Reingold, and Richard Zemel. 2012.
\newblock Fairness through awareness.
\newblock In \emph{Proceedings of the 3rd innovations in theoretical computer science conference}, pages 214--226.

\bibitem[{Eickhoff(2014)}]{eickhoff2014crowd}
Carsten Eickhoff. 2014.
\newblock Crowd-powered experts: Helping surgeons interpret breast cancer images.
\newblock In \emph{Proceedings of the First International Workshop on Gamification for Information Retrieval}, pages 53--56.

\bibitem[{Eickhoff(2018)}]{eickhoff2018cognitive}
Carsten Eickhoff. 2018.
\newblock Cognitive biases in crowdsourcing.
\newblock In \emph{Proceedings of the eleventh ACM international conference on web search and data mining}, pages 162--170.

\bibitem[{Eickhoff et~al.(2012)Eickhoff, Harris, de~Vries, and Srinivasan}]{eickhoff2012quality}
Carsten Eickhoff, Christopher~G Harris, Arjen~P de~Vries, and Padmini Srinivasan. 2012.
\newblock Quality through flow and immersion: gamifying crowdsourced relevance assessments.
\newblock In \emph{Proceedings of the 35th international ACM SIGIR conference on Research and development in information retrieval}, pages 871--880.

\bibitem[{Ellsberg(1961)}]{ellsberg1961risk}
Daniel Ellsberg. 1961.
\newblock Risk, ambiguity, and the savage axioms.
\newblock \emph{The quarterly journal of economics}, 75(4):643--669.

\bibitem[{Friedler et~al.(2014)Friedler, Scheidegger, and Venkatasubramanian}]{friedler2014certifying}
Sorelle Friedler, Carlos Scheidegger, and Suresh Venkatasubramanian. 2014.
\newblock Certifying and removing disparate impact.
\newblock \emph{arXiv preprint arXiv:1412.3756}.

\bibitem[{Fromreide et~al.(2014)Fromreide, Hovy, and S{\o}gaard}]{fromreide2014crowdsourcing}
Hege Fromreide, Dirk Hovy, and Anders S{\o}gaard. 2014.
\newblock Crowdsourcing and annotating ner for twitter\# drift.
\newblock In \emph{LREC}, pages 2544--2547.

\bibitem[{Green and Hu(2018)}]{green2018myth}
Ben Green and Lily Hu. 2018.
\newblock The myth in the methodology: Towards a recontextualization of fairness in machine learning.
\newblock In \emph{Proceedings of the machine learning: the debates workshop}.

\bibitem[{Hajian et~al.(2016)Hajian, Bonchi, and Castillo}]{hajian2016algorithmic}
Sara Hajian, Francesco Bonchi, and Carlos Castillo. 2016.
\newblock Algorithmic bias: From discrimination discovery to fairness-aware data mining.
\newblock In \emph{Proceedings of the 22nd ACM SIGKDD international conference on knowledge discovery and data mining}, pages 2125--2126.

\bibitem[{Hardt et~al.(2016)Hardt, Price, and Srebro}]{hardt2016equality}
Moritz Hardt, Eric Price, and Nati Srebro. 2016.
\newblock Equality of opportunity in supervised learning.
\newblock \emph{Advances in neural information processing systems}, 29.

\bibitem[{Havens et~al.(2022)Havens, Terras, Bach, and Alex}]{havens2022uncertainty}
Lucy Havens, Melissa Terras, Benjamin Bach, and Beatrice Alex. 2022.
\newblock Uncertainty and inclusivity in gender bias annotation: An annotation taxonomy and annotated datasets of british english text.
\newblock In \emph{Proceedings of the 4th Workshop on Gender Bias in Natural Language Processing (GeBNLP)}, pages 30--57.

\bibitem[{Hettiachchi et~al.(2021)Hettiachchi, Sanderson, Goncalves, Hosio, Kazai, Lease, Schaekermann, and Yilmaz}]{hettiachchi2021investigating}
Danula Hettiachchi, Mark Sanderson, Jorge Goncalves, Simo Hosio, Gabriella Kazai, Matthew Lease, Mike Schaekermann, and Emine Yilmaz. 2021.
\newblock Investigating and mitigating biases in crowdsourced data.
\newblock In \emph{Companion Publication of the 2021 Conference on Computer Supported Cooperative Work and Social Computing}, pages 331--334.

\bibitem[{Hovy and Prabhumoye(2021)}]{hovy2021five}
Dirk Hovy and Shrimai Prabhumoye. 2021.
\newblock Five sources of bias in natural language processing.
\newblock \emph{Language and linguistics compass}, 15(8):e12432.

\bibitem[{Irani and Silberman(2013)}]{irani2013turkopticon}
Lilly~C Irani and M~Six Silberman. 2013.
\newblock Turkopticon: Interrupting worker invisibility in amazon mechanical turk.
\newblock In \emph{Proceedings of the SIGCHI conference on human factors in computing systems}, pages 611--620.

\bibitem[{Joseph et~al.(2016)Joseph, Kearns, Morgenstern, and Roth}]{joseph2016fairness}
Matthew Joseph, Michael Kearns, Jamie~H Morgenstern, and Aaron Roth. 2016.
\newblock Fairness in learning: Classic and contextual bandits.
\newblock \emph{Advances in neural information processing systems}, 29.

\bibitem[{Kamiran and Calders(2012)}]{kamiran2012data}
Faisal Kamiran and Toon Calders. 2012.
\newblock Data preprocessing techniques for classification without discrimination.
\newblock \emph{Knowledge and information systems}, 33(1):1--33.

\bibitem[{Khowaja et~al.(2023)Khowaja, Khuwaja, and Dev}]{khowaja2023chatgpt}
Sunder~Ali Khowaja, Parus Khuwaja, and Kapal Dev. 2023.
\newblock Chatgpt needs spade (sustainability, privacy, digital divide, and ethics) evaluation: A review.
\newblock \emph{arXiv preprint arXiv:2305.03123}.

\bibitem[{Kleinberg et~al.(2016)Kleinberg, Mullainathan, and Raghavan}]{kleinberg2016inherent}
Jon Kleinberg, Sendhil Mullainathan, and Manish Raghavan. 2016.
\newblock Inherent trade-offs in the fair determination of risk scores.
\newblock \emph{arXiv preprint arXiv:1609.05807}.

\bibitem[{Krasanakis et~al.(2018)Krasanakis, Spyromitros-Xioufis, Papadopoulos, and Kompatsiaris}]{krasanakis2018adaptive}
Emmanouil Krasanakis, Eleftherios Spyromitros-Xioufis, Symeon Papadopoulos, and Yiannis Kompatsiaris. 2018.
\newblock Adaptive sensitive reweighting to mitigate bias in fairness-aware classification.
\newblock In \emph{Proceedings of the 2018 world wide web conference}, pages 853--862.

\bibitem[{Lee and Floridi(2021)}]{lee2021algorithmic}
Michelle Seng~Ah Lee and Luciano Floridi. 2021.
\newblock Algorithmic fairness in mortgage lending: from absolute conditions to relational trade-offs.
\newblock \emph{Minds and Machines}, 31(1):165--191.

\bibitem[{Lee and Singh(2021)}]{lee2021landscape}
Michelle Seng~Ah Lee and Jat Singh. 2021.
\newblock The landscape and gaps in open source fairness toolkits.
\newblock In \emph{Proceedings of the 2021 CHI conference on human factors in computing systems}, pages 1--13.

\bibitem[{Lewicki et~al.(2023)Lewicki, Lee, Cobbe, and Singh}]{lewicki2023out}
Kornel Lewicki, Michelle Seng~Ah Lee, Jennifer Cobbe, and Jatinder Singh. 2023.
\newblock Out of context: Investigating the bias and fairness concerns of “artificial intelligence as a service”.
\newblock In \emph{Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems}, pages 1--17.

\bibitem[{Liu et~al.(2017)Liu, Radanovic, Dimitrakakis, Mandal, and Parkes}]{liu2017calibrated}
Yang Liu, Goran Radanovic, Christos Dimitrakakis, Debmalya Mandal, and David~C Parkes. 2017.
\newblock Calibrated fairness in bandits.
\newblock \emph{arXiv preprint arXiv:1707.01875}.

\bibitem[{Martie et~al.(2005)Martie, Nettle, and Murray}]{martie2005evolution}
Haselton~G Martie, Daniel Nettle, and Damian~R Murray. 2005.
\newblock The evolution of cognitive bias. 724-746 in the handbook of evolutionary psychology, edited by david m. buss.

\bibitem[{Mehrabi et~al.(2021)Mehrabi, Morstatter, Saxena, Lerman, and Galstyan}]{mehrabi2021survey}
Ninareh Mehrabi, Fred Morstatter, Nripsuta Saxena, Kristina Lerman, and Aram Galstyan. 2021.
\newblock A survey on bias and fairness in machine learning.
\newblock \emph{ACM computing surveys (CSUR)}, 54(6):1--35.

\bibitem[{Parmar et~al.(2022)Parmar, Mishra, Geva, and Baral}]{parmar2022don}
Mihir Parmar, Swaroop Mishra, Mor Geva, and Chitta Baral. 2022.
\newblock Don't blame the annotator: Bias already starts in the annotation instructions.
\newblock \emph{arXiv preprint arXiv:2205.00415}.

\bibitem[{Plank et~al.(2014)Plank, Hovy, and S{\o}gaard}]{plank2014linguistically}
Barbara Plank, Dirk Hovy, and Anders S{\o}gaard. 2014.
\newblock Linguistically debatable or just plain wrong?
\newblock In \emph{Proceedings of the 52nd Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers)}, pages 507--511.

\bibitem[{Pleiss et~al.(2017)Pleiss, Raghavan, Wu, Kleinberg, and Weinberger}]{pleiss2017fairness}
Geoff Pleiss, Manish Raghavan, Felix Wu, Jon Kleinberg, and Kilian~Q Weinberger. 2017.
\newblock On fairness and calibration.
\newblock \emph{Advances in neural information processing systems}, 30.

\bibitem[{Quinn and Bederson(2011)}]{quinn2011human}
Alexander~J Quinn and Benjamin~B Bederson. 2011.
\newblock Human computation: a survey and taxonomy of a growing field.
\newblock In \emph{Proceedings of the SIGCHI conference on human factors in computing systems}, pages 1403--1412.

\bibitem[{Reinecke and Gajos(2015)}]{reinecke2015labinthewild}
Katharina Reinecke and Krzysztof~Z Gajos. 2015.
\newblock Labinthewild: Conducting large-scale online experiments with uncompensated samples.
\newblock In \emph{Proceedings of the 18th ACM conference on computer supported cooperative work \& social computing}, pages 1364--1378.

\bibitem[{Rudolph et~al.(2023)Rudolph, Tan, and Tan}]{rudolph2023chatgpt}
J{\"u}rgen Rudolph, Samson Tan, and Shannon Tan. 2023.
\newblock Chatgpt: Bullshit spewer or the end of traditional assessments in higher education?
\newblock \emph{Journal of Applied Learning and Teaching}, 6(1).

\bibitem[{Sambasivan et~al.(2021)Sambasivan, Arnesen, Hutchinson, Doshi, and Prabhakaran}]{sambasivan2021re}
Nithya Sambasivan, Erin Arnesen, Ben Hutchinson, Tulsee Doshi, and Vinodkumar Prabhakaran. 2021.
\newblock Re-imagining algorithmic fairness in india and beyond.
\newblock In \emph{Proceedings of the 2021 ACM conference on fairness, accountability, and transparency}, pages 315--328.

\bibitem[{Sap et~al.(2019)Sap, Card, Gabriel, Choi, and Smith}]{sap2019risk}
Maarten Sap, Dallas Card, Saadia Gabriel, Yejin Choi, and Noah~A Smith. 2019.
\newblock The risk of racial bias in hate speech detection.
\newblock In \emph{Proceedings of the 57th annual meeting of the association for computational linguistics}, pages 1668--1678.

\bibitem[{Seaborn and Kim(2023)}]{seaborn2023m}
Katie Seaborn and Yeongdae Kim. 2023.
\newblock “i'm” lost in translation: Pronoun missteps in crowdsourced data sets.
\newblock In \emph{Extended Abstracts of the 2023 CHI Conference on Human Factors in Computing Systems}, pages 1--6.

\bibitem[{Selbst et~al.(2019)Selbst, Boyd, Friedler, Venkatasubramanian, and Vertesi}]{selbst2019fairness}
Andrew~D Selbst, Danah Boyd, Sorelle~A Friedler, Suresh Venkatasubramanian, and Janet Vertesi. 2019.
\newblock Fairness and abstraction in sociotechnical systems.
\newblock In \emph{Proceedings of the conference on fairness, accountability, and transparency}, pages 59--68.

\bibitem[{Shrestha et~al.(2022)Shrestha, Kafle, and Kanan}]{shrestha2022investigation}
Robik Shrestha, Kushal Kafle, and Christopher Kanan. 2022.
\newblock An investigation of critical issues in bias mitigation techniques.
\newblock In \emph{Proceedings of the IEEE/CVF Winter Conference on Applications of Computer Vision}, pages 1943--1954.

\bibitem[{Simon(1957)}]{simon1957models}
Herbert~A Simon. 1957.
\newblock Models of man; social and rational.

\bibitem[{Sun et~al.(2019)Sun, Gaut, Tang, Huang, ElSherief, Zhao, Mirza, Belding, Chang, and Wang}]{sun2019mitigating}
Tony Sun, Andrew Gaut, Shirlyn Tang, Yuxin Huang, Mai ElSherief, Jieyu Zhao, Diba Mirza, Elizabeth Belding, Kai-Wei Chang, and William~Yang Wang. 2019.
\newblock Mitigating gender bias in natural language processing: Literature review.
\newblock \emph{arXiv preprint arXiv:1906.08976}.

\bibitem[{Suresh and Guttag(2021)}]{suresh2021framework}
Harini Suresh and John Guttag. 2021.
\newblock A framework for understanding sources of harm throughout the machine learning life cycle.
\newblock In \emph{Equity and access in algorithms, mechanisms, and optimization}, pages 1--9.

\bibitem[{Talboy and Schneider(2022)}]{talboy2022reference}
Alaina Talboy and Sandra Schneider. 2022.
\newblock Reference dependence in bayesian reasoning: Value selection bias, congruence effects, and response prompt sensitivity.
\newblock \emph{Frontiers in Psychology}, 13:729285.

\bibitem[{Thoppilan et~al.(2022)Thoppilan, De~Freitas, Hall, Shazeer, Kulshreshtha, Cheng, Jin, Bos, Baker, Du et~al.}]{thoppilan2022lamda}
Romal Thoppilan, Daniel De~Freitas, Jamie Hall, Noam Shazeer, Apoorv Kulshreshtha, Heng-Tze Cheng, Alicia Jin, Taylor Bos, Leslie Baker, Yu~Du, et~al. 2022.
\newblock Lamda: Language models for dialog applications.
\newblock \emph{arXiv preprint arXiv:2201.08239}.

\bibitem[{Tversky and Kahneman(1974)}]{tversky1974judgment}
Amos Tversky and Daniel Kahneman. 1974.
\newblock Judgment under uncertainty: Heuristics and biases: Biases in judgments reveal some heuristics of thinking under uncertainty.
\newblock \emph{science}, 185(4157):1124--1131.

\bibitem[{Van~Berkel et~al.(2019)Van~Berkel, Goncalves, Hettiachchi, Wijenayake, Kelly, and Kostakos}]{van2019crowdsourcing}
Niels Van~Berkel, Jorge Goncalves, Danula Hettiachchi, Senuri Wijenayake, Ryan~M Kelly, and Vassilis Kostakos. 2019.
\newblock Crowdsourcing perceptions of fair predictors for machine learning: A recidivism case study.
\newblock \emph{Proceedings of the ACM on Human-Computer Interaction}, 3(CSCW):1--21.

\bibitem[{Van~Berkel et~al.(2021)Van~Berkel, Goncalves, Russo, Hosio, and Skov}]{van2021effect}
Niels Van~Berkel, Jorge Goncalves, Daniel Russo, Simo Hosio, and Mikael~B Skov. 2021.
\newblock Effect of information presentation on fairness perceptions of machine learning predictors.
\newblock In \emph{Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems}, pages 1--13.

\bibitem[{Van~Dis et~al.(2023)Van~Dis, Bollen, Zuidema, van Rooij, and Bockting}]{van2023chatgpt}
Eva~AM Van~Dis, Johan Bollen, Willem Zuidema, Robert van Rooij, and Claudi~L Bockting. 2023.
\newblock Chatgpt: five priorities for research.
\newblock \emph{Nature}, 614(7947):224--226.

\bibitem[{Verma and Rubin(2018)}]{verma2018fairness}
Sahil Verma and Julia Rubin. 2018.
\newblock Fairness definitions explained.
\newblock In \emph{Proceedings of the international workshop on software fairness}, pages 1--7.

\bibitem[{Wang et~al.(2013)Wang, Wang, Li, Zheng, and Zhao}]{wang2013characterizing}
Tianyi Wang, Gang Wang, Xing Li, Haitao Zheng, and Ben~Y Zhao. 2013.
\newblock Characterizing and detecting malicious crowdsourcing.
\newblock In \emph{Proceedings of the ACM SIGCOMM 2013 conference on SIGCOMM}, pages 537--538.

\bibitem[{Wu et~al.(2022)Wu, Xiao, Sun, Zhang, Ma, and He}]{wu2022survey}
Xingjiao Wu, Luwei Xiao, Yixuan Sun, Junhang Zhang, Tianlong Ma, and Liang He. 2022.
\newblock A survey of human-in-the-loop for machine learning.
\newblock \emph{Future Generation Computer Systems}, 135:364--381.

\bibitem[{Zafar et~al.(2017)Zafar, Valera, Gomez~Rodriguez, and Gummadi}]{zafar2017fairness}
Muhammad~Bilal Zafar, Isabel Valera, Manuel Gomez~Rodriguez, and Krishna~P Gummadi. 2017.
\newblock Fairness beyond disparate treatment \& disparate impact: Learning classification without disparate mistreatment.
\newblock In \emph{Proceedings of the 26th international conference on world wide web}, pages 1171--1180.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage[final]{acl}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\title{Blind Spots and Biases: Exploring the Role of Annotator Cognitive Biases in NLP}

\author{Sanjana Gautam\\
    Information Sciences and Technology \\
    Pennsylvania State University, USA\\
  \texttt{sanjana.gautam@psu.edu}\\
    \And
    Mukund Srinath \\
    Information Sciences and Technology \\
    Pennsylvania State University, USA\\
  \texttt{mukund@psu.edu}}

\begin{document}
\maketitle
\begin{abstract}
With the rapid proliferation of artificial intelligence, there is growing concern over its potential to exacerbate existing biases and societal disparities and introduce novel ones. This issue has prompted widespread attention from academia, policymakers, industry, and civil society. While evidence suggests that integrating human perspectives can mitigate bias-related issues in AI systems, it also introduces challenges associated with cognitive biases inherent in human decision-making. Our research focuses on reviewing existing methodologies and ongoing investigations aimed at understanding annotation attributes that contribute to bias.
\end{abstract}

\section{Introduction}

With the recent rapid expansion of generative AI models, we have witnessed their numerous benefits and the emergence of substantial ethical concerns \cite{thoppilan2022lamda, rudolph2023chatgpt}. There has been an influx of remarkable and noteworthy work that describes the issues of fairness, toxicity and bias in the text generation process \cite{bender2021dangers, abid2021persistent, seaborn2023m}. Several models are deployed as real-world solutions with a lack of informed consideration of their social implications, especially in sensitive fields such as healthcare, journalism, law, and finance \cite{khowaja2023chatgpt}. Recent research has revealed that these language models can mimic human biases present in language, perpetuating prejudiced behaviour that dehumanizes certain socio-demographic groups by deeming them more negative or toxic \cite{havens2022uncertainty, blodgett2020language}. 

One of the proposed solutions to this issue has been to introduce human annotators to label the training corpora or validate pre-labelled datasets and manually remove toxic (or biased) data entries \cite{havens2022uncertainty, cabrera2014systematic}. It is common practice for machine learning systems to rely on crowd-sourced label data for training and evaluation \cite{wu2022survey}. It is also well-known that biases present in the label data can induce biases in the trained models \cite{hettiachchi2021investigating}. Therefore, while humans-in-the-loop for model training may seem like an intuitive solution, it often introduces additional biases due to inherent cognitive biases in humans \citep{parmar2022don}. Crowdwork annotation studies conducted on MTurk (and other crowdwork platforms) where the participants come from a specific demographic population can potentially perpetuate populist viewpoints \citep{reinecke2015labinthewild}. 

Prior work has established the pitfalls in human rationality, as influenced by the lived experiences and environment, which Herbert Simon termed bounded rationality \cite{simon1957models}. Human biases have been identified to be the resulting gap between rational behaviour and heuristically determined behaviour \cite{tversky1974judgment, bojke2021developing}. Over 180 cognitive biases have been identified, spawning everything from social interaction to judgment and decision-making with research spanning over 70 years \cite{talboy2022reference}. These tendencies or patterns can lead to faulty reasoning, irrationality, and potentially detrimental outcomes.

Bias sometimes emerges due to distractions, lack of interest, or laziness among annotators regarding the annotation task, leading them to select inaccurate labels. However, more concerning is the label bias stemming from informed and well-intentioned annotators who consistently exhibit disagreement \cite{hovy2021five}. \citet{plank2014linguistically} demonstrated that this form of bias emerges when multiple correct labels are possible. For instance, the term 'social media' can be legitimately interpreted either as a noun phrase consisting of an adjective and a noun or as a noun compound comprising two nouns. For example, \citet{sap2019risk} demonstrated that these biases mirror social and demographic variances. For instance, annotators tend to evaluate utterances from various ethnic groups disparately and may misinterpret harmless banter as hate speech due to their unfamiliarity with the communication norms of the original speakers.

Merely relying on a few gold-standard corpora as training datasets or debiasing datasets is not a sustainable long-term strategy since languages undergo constant evolution. Thus, even a comprehensive sample can only encapsulate a momentary snapshot, offering at best a transient solution \cite{fromreide2014crowdsourcing}. We believe the design and set-up of the crowd work task plays a pivotal role in determining the goodness of data. In this work, we look at bias-diminishing strategies and identify the pressing questions in this area. Our central goal is to show that there is a need for standardized design principles when it comes to designing crowdwork studies. Specifically, we concentrate on the need for an HCI perspective in natural language processing research.

\section{Bias in AI Models}

Generative AI's propensity to amplify existing biases and create new ones has attracted considerable attention across a range of communities, including academics, policy-makers, industry, and civil society. Much of the initial work focused on developing quantitative definitions of fairness  \cite{dwork2012fairness, hardt2016equality, joseph2016fairness, liu2017calibrated, verma2018fairness}, and various technical methods for ‘debiasing’ AI models \cite{agarwal2018reductions, bolukbasi2016man, friedler2014certifying, zafar2017fairness}. When referring to de-biasing, we use the definition \textit{'removing undesired skews in the data and the model outcome, such as by equalising a metric of interest between groups'}. \textit{“Unintended bias”} is used to describe the different sources of bias that are introduced throughout an AI development life cycle \cite{lee2021algorithmic, suresh2021framework}, focusing on not just the bias introduced, but also the harm it causes \cite{crawford2016artificial}.

Recent studies have shifted focus from merely identifying sources of bias in AI, such as flawed data collection methods, to exploring the various harms caused by these biases. This shift is supported by interdisciplinary research that highlights the contextual nature of fairness. Factors such as regional and cultural differences in lived experiences significantly influence perceptions of fairness, revealing that certain algorithmic behaviours may only be deemed harmful in specific social or cultural contexts \cite{green2018myth, lee2021landscape, sambasivan2021re, selbst2019fairness}. Given these complexities, it is broadly acknowledged that eliminating bias or ensuring absolute fairness in AI systems is unfeasible \cite{kleinberg2016inherent, mehrabi2021survey, pleiss2017fairness}. Instead, the objective is to minimize fairness-related harms and other adverse impacts to the greatest extent possible \cite{mehrabi2021survey, selbst2019fairness, sun2019mitigating}. This perspective is further enhanced by recent interdisciplinary studies \cite{lewicki2023out}, which underscore the nuanced and multifaceted nature of fairness in AI.

Identifying and acknowledging systemic biases in data collection is a crucial step in mitigating their impact on the systems that are trained using this data, and is a critical prerequisite for achieving fairness in algorithmic decision-making \cite{hajian2016algorithmic}. While humans are integral to the system, participating in data collection and various phases thereof, it is imperative to emphasize that human computation \cite{quinn2011human}, the practice of harnessing human intelligence and cognitive abilities as computational elements, holds potential for addressing and mitigating these challenges.

\section{Cognitive Bias among Annotators}


As emphasized by \citet{van2023chatgpt}, ensuring human accountability is essential in scientific practice. The history of Large Language Models (LLMs) has shown that they can produce inaccurate information, or "hallucinations." To guarantee the accuracy of information, it is necessary to implement a rigorous verification and fact-checking process led by experts. Consequently, the discourse highlights the critical need for accountability in human-in-the-loop systems, particularly in response to the new challenges posed by these systems.

The importance of understanding and mitigating biases in crowd data is highly relevant to researchers, and others who rely on crowd data for creating automated systems. Prior work has explored various approaches to promoting fairness in machine learning, including the direct utilization of crowdsourced data \cite{balayn2018characterising}, leveraging crowds to assess perceived fairness of features \cite{van2019crowdsourcing, van2021effect}, applying pre-processing techniques such as removing sensitive attributes, resampling data to remove discrimination, and iteratively adjusting training weights for sensitive groups \cite{calmon2017optimized, kamiran2012data, krasanakis2018adaptive}, as well as employing active learning methods \cite{anahideh2022fair}.

The use of crowdsourcing for tasks such as data annotation can inadvertently introduce cognitive biases, stemming from the inherent design of the task itself. We have identified three primary reasons why annotated data can be problematic: (1) Unethical spammers submit imprecise or even arbitrary labels in order to maximize their financial advantage \cite{eickhoff2012quality} or due to external distractions. (2) Unqualified workers are, despite their best efforts, unable to produce an acceptable annotation quality \cite{eickhoff2014crowd}. (3) Malicious workers purposefully aim to undermine or influence the labelling effort \cite{wang2013characterizing}. However, we propose that there might be some factors that have not been uncovered in prior literature. Crowd-workers have their tasks cut out for them, in cases where the nature of task design causes the propagation of bias. Research on crowd work has often focused on task accuracy whereas other factors such as biases in data have received limited attention \cite{hettiachchi2021investigating}. 

Cognitive biases originate from individuals' own \textit{"subjective social reality"} which is often a product of lived experiences. This makes cognitive bias a deviation from the rationality of judgement, therefore it may consist of perceptions of other people that are often illogical \cite{martie2005evolution}.  An individual’s construction of social reality, instead of the objective input, may dictate their behaviour and lead to perceptual distortion, inaccurate judgment, illogical interpretation, or irrationality \cite{bless2014social}. Past work has demonstrated that cognitive bias can affect crowdsourced labour and lead to significantly reduced result quality. This performance detriment is subsequently propagated into system ranking robustness and machine-learned ranker efficacy  \cite{eickhoff2018cognitive}.


The annotation instructions provided to crowdworkers can inadvertently prime them to exhibit biases towards or against specific domain information, which can be exacerbated by poorly designed instructions. Furthermore, annotators are often not fully informed about the true purpose of the research, leading to an ambiguity effect that can make the decision-making process appear more challenging and less appealing due to the limited information available \cite{ellsberg1961risk}. Additionally, the phased revelation of information to annotators can result in an anchoring effect, where certain pieces of information are given disproportionate attention based on the timing of their disclosure. This underscores the importance of designing annotation studies that mitigate cognitive biases among workers, ensuring that the annotation process is fair, transparent, and unbiased.

\section{Crowd Control}

Humans in the loop bring a lot of value to generative AI and AI systems. Therefore, the solution to the issue of cognitive bias cannot be to remove the annotators from the system. Human annotators often bring expert judgements, that are valuable in creating ground truth labels. For example, annotation of medical imagery cannot be performed without the help of annotators who are medical professionals. Expert guidance, lived experiences and proximity to the problem domain make human annotators irreplaceable in the AI-training life-cycle. The common strategies of accounting for biases of annotators by employing qualification tests, demographic filters, incentives, and sophisticated worker models may not be enough to overcome this source of noise. There is therefore a need to control the annotation task design settings, to minimize the introduction of biases due to the cognitive biases of annotators. 
While cognitive biases and their effects on decision-making are well-known and widely studied, we note that AI-assisted decision-making presents a new decision-making paradigm. It is important to study their role in this new paradigm, both analytically and empirically.

Crowdwork platforms are often designed to position crowdworkers as interchangeable \cite{irani2013turkopticon}. While some forms of digital work can be decomposed and distributed, the presumption that all crowdsourced dataset annotators exercise near-identical capacities of perception and judgement ignores the fact that social position, identity, and experience shape how annotators' actions.

Previous research has highlighted the significance of the annotator population and the power dynamics inherent in platform-mediated crowdwork, both of which can perpetuate cognitive biases \cite{diaz2022crowdworksheets}. Building upon this foundation, we propose a novel framework to enhance transparency and robustness in the process of designing a crowdwork task. This approach holds promise for mitigating the impact of cognitive biases in crowdwork, thereby contributing to more reliable and trustworthy outcomes.

\section{Counter-measures for Biases}



To minimize bias in NLP annotation tasks, several steps can be implemented. Firstly, recruiting a diverse group of annotators from various backgrounds can help balance individual biases. Providing clear and detailed guidelines ensures uniform understanding across annotators. Training sessions, followed by calibration discussions, align annotator interpretations and reveal guideline ambiguities. An iterative feedback loop allows for regular quality checks and guideline adjustments based on annotator experiences. Measuring inter-annotator agreement with metrics like Cohen's Kappa highlights discrepancies and areas needing clarification. Annotation tasks should be designed to minimize bias, such as by rotating text assignments among annotators to avoid topical biases. Finally, a post-annotation analysis can detect any remaining biases, ensuring the reliability and fairness of the annotated data. 

However, biases can arise at any point in the AI lifecycle. It is therefore imperative for researchers to maintain a meticulous approach throughout the entire research process, encompassing various facets such as the selection of appropriate datasets, adherence to annotation schemes or labelling procedures, thoughtful considerations regarding data representation methodologies, judicious selection of algorithms tailored to the task at hand, and rigorous evaluation protocols for automated systems. Furthermore, researchers must consider the tangible real-world applications of their research endeavors. Particularly noteworthy is the imperative to consciously direct efforts towards leveraging technological advancements to uplift and empower marginalized communities, as underscored by \citet{asad2019academic}. Several studies critique existing bias mitigation algorithms for their lack of effectiveness due to inconsistent study protocols, inappropriate datasets, and over-tuning to specific test sets. To overcome these limitations, research needs to introduce robust evaluation protocol, and sensible metrics designed to evaluate algorithm robustness against various biases \citep{shrestha2022investigation}. 

Our future work derives from the insights presented in the preceding discussion. It posits that the roots of bias within AI systems often traced back to the initial stages of the annotation process, particularly during the instruction phase. Although not all cognitive biases are inherently detrimental, a pressing need exists to advance our comprehension of how to devise annotation studies that align with the principles of human-computer interaction (HCI). 

Our objective in this research endeavour is to contribute substantively to the ongoing efforts aimed at mitigating bias in crowd work. We intend to achieve this by focusing on the refinement of study design and instructional strategies. By incorporating insights from the HCI discipline, we aim to cultivate a nuanced understanding of how to create balanced annotation studies that minimize the emergence of bias. Through this work, we aspire to not only shed light on the pivotal role played by the annotation phase in propagating or mitigating bias but also to provide practical recommendations and guidelines for researchers and practitioners engaged in AI development and crowd work.

\section{Conclusion}

Our research highlights the critical importance of considering annotation attributes that contribute to bias in AI systems. The cognitive biases of annotators, inherent in human decision-making, can perpetuate and even amplify existing social disparities in AI models. To mitigate these issues, a multidisciplinary approach is necessary not only in deploying AI models but also in designing better systems for annotation tasks. By bringing together experts from diverse fields, including human-centered design, ethics, social sciences, law, healthcare, AI/ML, education, communication, and community representation, we can design annotation systems that are more inclusive, transparent, and fair. This collaborative framework is essential for developing annotation tasks that are free from biases, ambiguous, and unclear instructions, and that take into account the complexities of real-world data. Furthermore, a multidisciplinary approach is crucial for deploying AI models that are developed using these annotated data, ensuring that they are fair, transparent, and accountable. By acknowledging the limitations of human annotators and addressing them through a multidisciplinary approach, we can work towards a more equitable digital landscape where AI systems benefit both individuals and society as a whole.

\bibliography{custom}

\appendix

\end{document}